Hi,
If x= log of 2 to the base 15; y= log of 3 to the base 10 and z=log of 5 to the base 6 then evaluate xy+yz+zx+2xyz
Regards,
AN

Answer by lwsshak3(11628)

If x= log of 2 to the base 15; y= log of 3 to the base 10 and z=log of 5 to the base 6 then evaluate xy+yz+zx+2xyz
**
Convert to exponential form: (base raised to log of number=number)
15^x=2
10^y=3
6^z=5
..
xlog15=log2
x=log2/log15
x=.2560
..
ylog10=log3
y=log3/log10
y=.4771
..
zlog6=log5
z=log5/log6
z=.8982
..
xy+yz+zx+2xyz
=.2560*.4771+.4771*.8982+.8982*2360+2*.2560*.4771*.8982
=.1221+.4285+.2120+.2194
=.9820